OpenStudy (anonymous):

Given the equations 9x+34y=6 and 2x+12y=9, by what factor would you multiply the second equation to eliminate y and solve the system through linear combinations?

OpenStudy (jdoe0001):

well.. you want to eliminate "y" so your two "y's" are 34y and 12y by what factor do you multiply the 2nd "y", or 12y to eliminate "y",or to get a 0 well... call the factor.... "a" we know then that, 12y * a when added to 34y gives 0 so 12y * a + 34y = 0 solve for "a", see what you get

OpenStudy (anonymous):

Thank you I will do so :)

OpenStudy (anonymous):

Hmmm.. the answers contain a fraction and I am not getting a fraction in the multiple choice.

OpenStudy (whpalmer4):

how about you verify that you typed the equations correctly, and tell us what the answer choices are?

OpenStudy (anonymous):

Well, yes I did type them correctly and the choices are: -4/3 -3/4 -3/2 and -7/2

OpenStudy (whpalmer4):

Sorry, I apparently misread your response as "there aren't any fractions in the answer choices" but should have read it as "I get a fraction which isn't in the answer choices". I agree that the fraction I get is not on the list of choices.

OpenStudy (anonymous):

Yeah this math is so strange

OpenStudy (whpalmer4):

Well, mostly strange because somewhere along the line someone made an error or the problem was otherwise corrupted. Always frustrating to try to understand something new when the materials are incorrect!

OpenStudy (anonymous):

Yeah, I don't enjoy this kind of math and if I don't go through it I will fail my senior year :/. I am in desperate need of passing.

OpenStudy (whpalmer4):

It isn't so bad once you get the hang of it. You can always eliminate one equation by multiplying each equation by the opposite equation's coefficient of the variable you are trying to eliminate. For example: \[2x + 3y = 6\]\[3x + 4y=1\]If you want to eliminate \(y\), multiply the first equation by 4, and the second equation by -3 (I changed the sign to make one result be negative). \[4*2x + 4*3y = 4*6\]\[-3*3x-3*4y=-3*1\]or\[8x+12y=24\]\[-9x-12y=-3\] Notice now that we have equal but opposite coefficients for the variable we want to eliminate, and if we add the two equations together, we get \[8x-9x+12y-12y=24-3\]\[-x+0y=21\]\[-x=21\]\[x=-21\] Plugging that into one of the original formulas\[2(-21)+3y=6\]\[-42+3y=6\]\[3y=48\]\[y=16\] So our solution is \((-21,16)\) That wasn't so bad, was it?

OpenStudy (anonymous):

Nah, but when you get a defective program that doesn't teach you the correct things. It's hard to work around. I understand better and will take notes though for the test. Thank you.

OpenStudy (whpalmer4):

Good luck, I hope my explanation helped!